# VGS-ATD: Robust Distributed Learning for Multi-Label Medical Image Classification Under Heterogeneous and Imbalanced Conditions

### v1

@@ III Methodology
In this section, we provide a detailed introduction to the proposed VGS-ATD, including its overall workflow, methodology, and specific design for its three configurations, which are tailored to different scenarios.
@@
In the following subsections, we present multiple configurations of the VGS-ATD, each tailored to address different practical medical scenarios. An illustrative overview of these configurations is provided in Fig. 3.
Figure 3: (a), A sample of the VGS-ATD’s horizontal training configuration when two nodes share the same feature space. (b), A sample of the VGS-ATD’s vertical training configuration when two nodes do not have overlap in feature space. (c), A sample of the VGS-ATD’s ATD in ATD configuration.
III-A Configuration of Horizontal VGS-ATD
This is the basic configuration of the VGS-ATD, which aims to expand the training dataset across client nodes without sharing raw data by using a decentralized paradigm. The client nodes under this configuration share the same feature space or have overlapping feature distributions. Specifically, we conduct a node-to-node knowledge transfer learning between client nodes to enable a better local model performance and avoid domain mismatch. Specifically, an initial client node $n_{1}$ trains a local ViT encoder $f^{1}_{\theta}$ on its private dataset $D^{(1)}={(x_{1},y_{1})}$ by minimizing the standard cross-entropy loss:
@@ III-A Configuration of Horizontal VGS-ATD
III-B Configuration of Vertical VGS-ATD
The second configuration targets scenarios where client nodes have non-overlapping feature spaces, such as data from different modalities or backgrounds. This setup enables the system to generalize across a broader range of medical conditions within a unified model. However, transferring a model from one node to another in such heterogeneous settings can cause domain mismatch, often leading to degraded performance. This limitation occurs with many pretraining methods, particularly when the source domain differs significantly from the target domain. For example, models pretrained on general RGB datasets, such as ImageNet, often underperform on grayscale medical images [21]. To address this, one designated client (with sufficient computational resources) trains a model on a large-scale, publicly available medical dataset (e.g., chest X-rays). The resulting backbone extractor weights $W_{p}$ are then distributed to all nodes as a shared foundation. Nodes may optionally train from scratch if the pretrained weights are not suitable.
@@ III-B Configuration of Vertical VGS-ATD
III-C Configuration of Hierarchical ATD in ATD
The last configuration, referred to as ATD-in-ATD, extends the previous two configurations by introducing a hierarchical model aggregation strategy to enhance system scalability and flexibility at minimal cost. The aim of developing this setting is to enable the integration of different node groups without discarding their previously learned knowledge. The process begins by grouping client nodes and aggregating their local models into intermediate global models using either vertical or horizontal learning. These intermediate global models are then treated as new base models and are further aggregated across groups. This hierarchical fusion strategy enables the ATD system to flexibly integrate previously trained models, allowing new incoming nodes to immediately contribute to or benefit from the collective knowledge without requiring retraining from scratch.
@@ IV-A Environment Setup
In our setup, each client updates its local model for 50 training epochs. The optimizer used is AdamW, with a learning rate set to 3e-3 and a weight decay of 0.05. AdamW is chosen based on findings from the original ViT paper [8], which compared its performance to SGD and vanilla Adam. Models trained with AdamW demonstrated better generalization and were more effective at mitigating overfitting, which is an important consideration in the medical imaging domain, as training data is often limited and ViTs are known to be data-hungry.
The learning rate was selected based on initial empirical observations. For example, when using a relatively large learning rate (e.g., 0.1), the model showed performance improvements during the first five epochs but then experienced a decline, suggesting overfitting due to large gradient steps. Conversely, setting the learning rate too low resulted in gradual improvement, but the model failed to converge to optimal performance within the 50-epoch limit. To ensure fair comparisons, all methods in the subsequent experiments are trained under the same hyperparameter settings.
For the downstream classifier hyperparameter settings, we evaluated various configurations for each downstream classifier on the test datasets and selected the settings that yielded the best performance. Specifically, the MLP classifier was trained using the Adam optimiser for 50 epochs with a learning rate of 0.001. The SVM classifier employed the squared_hinge loss function with a convergence tolerance of 0.01. For the Random Forest classifier, we used the Gini impurity as the splitting criterion, set max_features to log2, and specified min_samples_split as 3.
IV-B Dataset and Preprocessing
@@
For the horizontal configuration test, we selected three chest X-ray datasets that focus on lung diseases and exhibit similar feature distributions: Tuberculosis Chest X-ray Dataset[22], Pneumonia Chest X-ray Dataset[23], and Covid Chest X-ray Dataset[24].
@@
For the vertical configuration test, we expanded the modality coverage from chest X-rays to include four major grayscale medical imaging types: MRI, CT, X-ray, and ultrasound. This configuration incorporates 12 additional datasets spanning various anatomical regions and diagnostic tasks, including Fracture Spine Dataset, Fracture Body Dataset, Chest Covid CT Dataset[25], Kidney Cancer Dataset[26], Brain Tumor Dataset, PCOS Ultrasound Dataset, Liver CT Dataset, Brain Tumor CT Dataset, Liver Fibrosis Ultrasound Dataset[27], Breast Cancer MRI Dataset, Lumbar Coordinate Dataset, and Breast Tumor MRI Dataset. A total of 15 datasets are selected and stored in the corresponding simulated client nodes in this phase. This setting significantly increases the diversity and complexity of the target domain, enabling a robust evaluation of the method’s generalisability under modality shifts and data imbalance.
@@
Finally, for the hierarchical ATD-in-ATD configuration test, we introduce 14 color image datasets drawn from the domains of pathology, endoscopy, fundus imaging, and general clinical imaging. The aggregated global model of the color image group would then be combined with the model that we obtained from the previous horizontal configuration (which was trained with grey-scale images). The color image dataset group include the Acute Lymphoblastic Leukemia Dataset[28], Breast Cancer Histopathological Dataset[29], Cervical Cancer Pathology Dataset, Lung and Colon Cancer Dataset[30], Lymphoma Cancer Dataset[31], Oral Cancer Dataset, Melanoma Cancer Dataset, Blood Cell Dataset, WCE Curated Colon Disease Dataset[32], Diabetic Foot Dataset [33], Retinal Cataract Dataset, Prostate Pathology Dataset, Otoscopic Dataset, and ORIGA Retinal Glaucoma Dataset. The number of selected datasets further increased to 29 in this phase. These datasets differ widely in image appearance, disease types, and acquisition modalities, providing an extreme-case scenario for evaluating the method’s resilience to distribution shifts, multi-label classification, and class imbalance.
@@
Fig. 4 presents a comprehensive overview of all datasets used in this study. Subfigure (a) provides a hierarchical summary of each dataset, organized into three levels: the general anatomical region, the specific body area where the disease occurs, and detailed dataset information including the imaging modality and reference source. Subfigure (b) illustrates the distribution of colour versus grayscale images, revealing a nearly balanced composition between the two types (approximately 50:50). Subfigure (c) highlights that the distribution of most grey-scale modalities (X-ray, CT, and MRI) is fairly balanced, with each accounting for around 13% of the total dataset. In contrast, within the colour image datasets, 41.3% of samples are derived from pathology images, while endoscopy and general/fundus images contribute 5.7% and 6.3%, respectively. Subfigure (d) complements Subfigure (a) by showing the proportional sample contribution of each dataset and its associated body region in relation to the entire dataset collection.
@@
All images were resized to 224×224 pixels to accommodate the input requirements of ViT models and to remain within the computational constraints of the available hardware. Furthermore, for datasets containing classes with fewer than 1,000 samples, data augmentation was applied to mitigate the risk of overfitting and enhance model generalizability. The augmentation techniques included random rotation (--30° to +30°), random horizontal flipping, zooming (scaling factor between 0.8× and 1.2×), gaussian blur ($\sigma$ = 0.1 to 2.0), and brightness adjustment (factor between 0.5 and 1.5). A custom augmentation pipeline randomly selected training samples and applied three out of these five transformations to construct the final augmented training set.
This diverse and stratified dataset preparation strategy is designed to rigorously assess the VGS-ATD’s adaptability, robustness, and effectiveness under real-world medical conditions. It also facilitates direct comparison with conventional approaches, such as centralized learning, federated learning, and swarm learning, in the presence of heterogeneous data distributions.
IV-C Evaluation Metrics
In this study, we assessed the performance of all trained classification models using four key evaluation metrics: accuracy, precision, sensitivity, and F1-score. While accuracy is a commonly used metric and is suitable when class distributions are relatively balanced, it may present misleading results in the presence of class imbalance. In contrast, the F1-score provides a more informative and robust measure under imbalanced conditions, as it considers both precision and recall. Given the highly imbalanced nature of the datasets used in this study, we place greater emphasis on the F1-score when comparing the VGS-ATD with baseline approaches to ensure a more equitable and meaningful evaluation.
The calculation of each metric was done as follows:
$Accuracy=\frac{TP+TN}{TP+FP+TN+FN}$ (13)
$Precision=\frac{TP}{TP+FP},~{}Sensitivity=\frac{TP}{TP+FN}$ (14)
$F1score=\frac{2TP}{FP+FN+2TP}$ (15)
Where TP is true positive, FP is false positive, TN is true negative, and FN is false negative.
IV-D Baselines Setup
@@
IV-D1 Centralized learning setup
@@
IV-D2 Federated learning setup
The federated learning configuration adopts a personalized training protocol, inspired by methods such as Federated Representation Learning [17] and FedBaBU [18]. In this setup, each client node shares only its backbone extractor weights during aggregation while maintaining its own classifier to accommodate differences in label space and data distribution. All federated clients are aligned with the VGS-ATD’s node configuration, and identical train/test splits are applied. Each client performs 10 local training epochs per round, with a total of five communication rounds, ensuring that all models are trained for 50 epochs to enable a fair comparison.
IV-D3 Swarm learning setup
The swarm learning setup mirrors the federated learning structure in terms of local training strategy. However, instead of centralized weight aggregation, each node sequentially passes its backbone extractor weights to the next node. This setup avoids classification head conflicts by training node-specific heads, making it suitable for environments with non-overlapping class labels.
It is important to note that all experiments were simulated on a local workstation, and network communication costs were not factored into the evaluation.
V Results and Analysis
@@
V-A Phase1: Classification on 3 Nodes
In the first phase, experiments are conducted using three datasets assigned to simulated client nodes. The results for this 3-node classification scenario are presented in Table. II. These three nodes each contain an independent chest X-ray dataset, with each containing a label related to normal chest and another specific chest disease label, collectively covering four unique class labels. Due to overlapping feature and label spaces in the three target nodes, this phase evaluates the performance of VGS-ATD using its horizontal configuration. To ensure a fair comparison, we apply transfer learning to all baseline methods using a shared source domain: the CXR Dataset instead of ImageNet, to minimise domain mismatch risk. The selected CXR dataset contains approximately 72,000 chest X-ray samples spanning four lung disease categories.
Following the initial transfer from the CXR dataset, VGS-ATD further performs node-to-node knowledge transfer, which is included in the horizontal configuration. However, such inter-node transfer is inherently incompatible with centralized and federated learning frameworks. In centralized learning, all datasets are aggregated and jointly trained, losing their individual identities. In federated learning, the communication structure is limited to node-to-server interactions, thereby precluding node-to-node transfer.
TABLE II: Phase1 Performance Comparison on 3 nodes.
Method Acc $\uparrow$ Pre $\uparrow$ Sensi $\uparrow$ F1 $\uparrow$ Central Learning 96.22% 96.27% 99.06% 95.69% Federated Learning 83.15% 80.94% 78.32% 77.72% Swarm Learning 80.31% 77.46% 70.09% 69.41% VGS-ATD (Ours) 97.33% 96.66% 97.24% 96.93% $\uparrow$: The higher the better.
@@ V-A Phase1: Classification on 3 Nodes
Figure 5: Phase1 training time and computational resource cost comparison.
@@
Additionally, Fig. 5 compares the training time and computational resource costs across VGS-ATD and baseline approaches. Both metrics are measured in minutes and are comparable since all methods were trained on the same hardware. Training time cost refers to the time from the start of training until the system is ready for prediction, while computational resource cost accounts for the total training time across all local models and global classifiers (as each minute spent on the same training device can be considered as a consumption of a certain amount of computational resource cost). For methods like centralized learning, which train a single model, the two metrics are effectively equivalent. In contrast, parallel approaches such as federated learning and VGS-ATD train local models concurrently, so their training time cost is determined by the node that takes the most time, while the overall computational cost reflects the total effort across all nodes.
@@
As shown in the figure, the proposed and federated methods achieve significantly lower training time thanks to their parallel design. Unlike centralized and swarm learning, they allow simultaneous training across nodes, which improves efficiency and scalability. For computational resource cost, centralized learning shows the best result, as it benefits from a simpler pipeline without requiring additional global classifier training.
V-B Phase2: Classification on 15 Nodes
TABLE III: Phase2 performance comparison on 15-node classification task.

### v3

@@ 3 Methodology
In the following subsections, we present three configurations of the VGS-ATD, each tailored to address different practical medical scenarios.
Figure 3: (a), A sample of the VGS-ATD’s horizontal training configuration when two nodes share the same feature space. Mn represents the corresponding model n, and Wn represents the weights transferred from model n. The final aggregated model is called Mag. (b), A sample of the VGS-ATD’s vertical training configuration when two nodes do not have overlap in feature space. Mp represents a shared pretrained model, and only the extractor part is transferred and aggregated in this setting. (c), A sample of the VGS-ATD’s ATD in ATD configuration. Model aggregation can be performed in multiple level.
3.1 Configuration of Horizontal VGS-ATD
This is the basic configuration of the VGS-ATD, which aims to expand the training dataset across client nodes without sharing raw data by using a decentralized paradigm. The workflow can be referenced in Fig. 3 (a). The goal of this configuration is to utilize a broader range of distributed data while minimizing the administrative burden associated with data access permissions. It is best suited for scenarios in which client nodes share a common feature space or have overlapping feature distributions. Specifically, we conduct a node-to-node knowledge transfer learning between client nodes to enable a better local model performance and avoid domain mismatch. Specifically, an initial client node $n_{1}$ trains a local ViT encoder $f^{1}_{\theta}$ on its private dataset $D^{(1)}={(x_{1},y_{1})}$ by minimizing the standard cross-entropy loss:
@@ 3.1 Configuration of Horizontal VGS-ATD
3.2 Configuration of Vertical VGS-ATD
The second configuration targets scenarios where client nodes have non-overlapping feature spaces, such as data from different modalities or backgrounds. As illustrated in Fig. 3 (b), the development of this configuration is to enable the system to generalize across a broader range of medical conditions within a unified model, thereby reducing the need to train separate models for each disease and minimizing computational redundancy. However, transferring a model from one node to another in such heterogeneous settings can cause domain mismatch, often leading to degraded performance. This limitation occurs with many pretraining methods, particularly when the source domain differs significantly from the target domain. For example, models pretrained on general RGB datasets, such as ImageNet, often underperform on grayscale medical images [wen2021rethinking]. To address this, one designated client (with sufficient computational resources) trains a model on a large-scale, publicly available medical dataset (e.g., chest X-rays). The resulting backbone extractor weights $W_{p}$ are then distributed to all nodes as a shared foundation. Nodes may optionally train from scratch if the pretrained weights are not suitable.
@@ 3.2 Configuration of Vertical VGS-ATD
3.3 Configuration of Hierarchical ATD in ATD
The last configuration, referred to as ATD-in-ATD (see Fig. 3 (c)), aims to extend the previous two configurations by introducing a hierarchical model aggregation strategy to enhance system scalability and flexibility at minimal cost. The aim of developing this setting is to enable the integration of different node groups without discarding their previously learned knowledge. The meaning of this configuration is to address the scalability challenge in decentralized learning setups, allowing for seamless model expansion while avoiding redundant training and preserving the accumulated knowledge base. This process begins by grouping client nodes and aggregating their local models into intermediate global models using either vertical or horizontal learning. These intermediate global models are then treated as new base models and are further aggregated across groups. This hierarchical fusion strategy enables the ATD system to flexibly integrate previously trained models, allowing new incoming nodes to immediately contribute to or benefit from the collective knowledge without requiring retraining from scratch.
@@ 4.1 Environment Setup
In our setup, each client updates its local model for 50 training epochs. The optimizer used is AdamW, with a learning rate set to 3e-3 and a weight decay of 0.05. AdamW is chosen based on findings from the original ViT paper [dosovitskiy2020image], which demonstrated better generalization and performance than SGD and vanilla Adam optimizer. The learning rate was selected based on initial empirical observations. For example, when using a relatively large learning rate (e.g., 0.1), the model showed performance improvements during the first five epochs but then experienced a decline, suggesting overfitting due to large gradient steps. Conversely, setting the learning rate too low resulted in gradual improvement, but the model failed to converge to optimal performance within the 50-epoch limit.
For the downstream classifier hyperparameter settings, we evaluated various configurations for each downstream classifier on the test datasets and selected the settings that yielded the best performance. Specifically, the MLP classifier was trained using the Adam optimizer for 50 epochs with a learning rate of 0.001. The SVM classifier employed the squared_hinge loss function with a convergence tolerance of 0.01. For the Random Forest classifier, we used the Gini impurity as the splitting criterion, set max_features to log2, and specified min_samples_split as 3.
4.2 Dataset and Preprocessing
@@
For the horizontal configuration test, we selected three chest X-ray datasets that focus on lung diseases and exhibit similar feature distributions: Tuberculosis Chest X-ray Dataset [rahman2020reliable], Pneumonia Chest X-ray Dataset [kermany2018identifying], and Covid Chest X-ray Dataset [shastri2022cheximagenet].
@@
For the vertical configuration test, we expanded the modality coverage from chest X-rays to include four major grayscale medical imaging types: MRI, CT, X-ray, and ultrasound. This configuration incorporates 12 additional datasets spanning various anatomical regions and diagnostic tasks, including Fracture Spine Dataset, Fracture Body Dataset, Chest Covid CT Dataset [maftouni2021robust], Kidney Cancer Dataset [islam2022vision], Brain Tumor Dataset, PCOS Ultrasound Dataset, Liver CT Dataset, Brain Tumor CT Dataset, Liver Fibrosis Ultrasound Dataset [joo2023classification], Breast Cancer MRI Dataset, Lumbar Coordinate Dataset, and Breast Tumor MRI Dataset. A total of 15 datasets are selected and stored in the corresponding simulated client nodes in this phase. This setting significantly increases the diversity and complexity of the target domain, enabling a robust evaluation of the method’s generalisability under modality shifts and data imbalance.
@@
Finally, for the hierarchical ATD-in-ATD configuration test, we introduce 14 color image datasets drawn from the domains of pathology, endoscopy, fundus imaging, and general clinical imaging. The aggregated global model of the color image group would then be combined with the model that we obtained from the previous horizontal configuration (which was trained with grey-scale images). The color image dataset group include the Acute Lymphoblastic Leukemia Dataset [ghaderzadeh2022fast], Breast Cancer Histopathological Dataset [spanhol2015dataset], Cervical Cancer Pathology Dataset, Lung and Colon Cancer Dataset [borkowski2019lung], Lymphoma Cancer Dataset [orlov2010automatic], Oral Cancer Dataset, Melanoma Cancer Dataset, Blood Cell Dataset, WCE Curated Colon Disease Dataset [pogorelov2017kvasir], Diabetic Foot Dataset [alzubaidi2020dfu_qutnet], Retinal Cataract Dataset, Prostate Pathology Dataset, Otoscopic Dataset, and ORIGA Retinal Glaucoma Dataset. The number of selected datasets further increased to 29 in this phase. These datasets differ widely in image appearance, disease types, and acquisition modalities, providing an extreme-case scenario for evaluating the method’s resilience to distribution shifts, multi-label classification, and class imbalance.
@@
Fig. 4 presents a comprehensive overview of all datasets used in this study. Subfigure (a) provides a hierarchical summary of each dataset, organized into three levels: the general anatomical region, the specific body area where the disease occurs, and detailed dataset information including the imaging modality and reference source. Subfigure (b) illustrates the distribution of color versus grayscale images, revealing a nearly balanced composition between the two groups (approximately 50:50). Subfigure (c) highlights that the distribution of most grey-scale modalities (X-ray, CT, and MRI) is fairly balanced, with each accounting for around 13% of the total dataset. In contrast, within the colour image datasets, 41.3% of samples are derived from pathology images, while endoscopy and general/fundus images contribute 5.7% and 6.3%, respectively. Subfigure (d) complements Subfigure (a) by showing the proportional sample contribution of each dataset and its associated body region in relation to the entire dataset collection.
@@
All images were resized to 224×224 pixels to accommodate the input requirements of ViT models and to remain within the computational constraints of the available hardware. Furthermore, for datasets containing classes with fewer than 1,000 samples, data augmentation was applied to mitigate the risk of overfitting and enhance model generalizability. The augmentation techniques included random rotation (-30° to +30°), random horizontal flipping, zooming (scaling factor between 0.8× and 1.2×), gaussian blur ($\sigma$ = 0.1 to 2.0), and brightness adjustment (factor between 0.5 and 1.5). A custom augmentation pipeline randomly selected training samples and applied three out of these five transformations to construct the final augmented training set.
4.3 Evaluation Metrics
In this study, we assessed the performance of all trained classification models using four key evaluation metrics: accuracy, precision, sensitivity, and F1-score. Accuracy measures the proportion of correctly classified samples (both positive and negative) out of the total number of samples. Precision evaluates how many of the samples predicted as positive are truly positive, and sensitivity measures the ability of the model to correctly identify actual positive cases. F1-score is the harmonic mean of precision and sensitivity, balancing the trade-off between the two. Together, these metrics provide a comprehensive evaluation of both the overall correctness of the models and their ability to correctly identify positive cases, which is especially important in medical applications.
4.4 Baselines Setup
@@
4.4.1 Centralized learning setup
@@
4.4.2 Federated learning setup
The federated learning configuration adopts a state-of-the-art personalized training protocol called federated learning based on global loss decomposition (FedLD) [zeng2024tackling], which applies a novel loss decomposition to address the data heterogeneity challenge. In this setup, all base model hyperparameter setups are aligned with the VGS-ATD’s node configuration, and identical train/test splits are applied. The pretrained weights used in VGS-ATD training are also applied to initialize the base FL model. Each client performs 10 local training epochs per round, with a total of five communication rounds, ensuring that all models are trained for 50 epochs to enable a fair comparison.
4.4.3 Swarm learning setup
The swarm learning setup follows a novel framework called non-independently and identically distributed swarm learning (Non-IID SL) [gao2022new], which is designed for decentralized heterogeneous data learning. The base model setup also mirrors those we applied for VGS-ATD and FedLD baseline, and each model is trained for 5 rounds with 10 epochs each round to ensure a fair comparison.
5 Results and Analysis
@@
5.1 Phase1: Classification on 3 Nodes
In the first phase, experiments are conducted using three datasets assigned to simulated client nodes. The results for this 3-node classification scenario are presented in Table. 2. The initial three nodes each contain an independent chest X-ray dataset, with each containing a label related to normal chest and another specific chest disease label, collectively covering four unique class labels. Due to overlapping feature and label spaces in the three target nodes, this phase evaluates the performance of VGS-ATD using its horizontal configuration. To ensure a fair comparison, we apply transfer learning to all baseline methods using a shared source domain: using Chest X-ray Dataset instead of ImageNet, to minimise domain mismatch risk. The selected source dataset contains approximately 72,000 chest X-ray samples spanning four lung disease categories.
Following the initial transfer from the source dataset, VGS-ATD further performs node-to-node knowledge transfer, which is included in the horizontal configuration. However, such inter-node transfer is inherently incompatible with centralized and federated learning frameworks. In centralized learning, all datasets are aggregated and jointly trained, losing their individual identities. In the federated learning baseline, the communication structure is limited to node-to-server interactions, thereby precluding node-to-node transfer.
Table 2: Phase1 Performance Comparison on 3 nodes.
Method Acc $\uparrow$ Pre $\uparrow$ Sensi $\uparrow$ F1 $\uparrow$ Central Learning 96.22% 96.27% 99.06% 95.69% FedLD 83.15% 80.94% 78.32% 77.72% NonIID-SL 80.31% 77.46% 70.09% 69.41% VGS-ATD (Ours) 97.33% 96.66% 97.24% 96.93% $\uparrow$: The higher the better.
@@ 5.1 Phase1: Classification on 3 Nodes
Figure 5: P1 training time and computational cost comparison.
@@
Additionally, Fig. 5 compares the training time and computational resource costs across VGS-ATD and baseline approaches. Both metrics are measured in minutes and are comparable since all methods were trained on the same hardware. Training time cost refers to the time from the start of training until the system is ready for prediction, while computational resource cost accounts for the total training time across all local models and global classifiers (as each minute spent on the same training device can be considered as a consumption of a certain amount of computational resource cost). For methods like centralized learning, which train a single model, the two metrics are effectively equivalent. In contrast, parallel approaches such as Fed-LD and VGS-ATD train local models concurrently, so their training time cost is determined by the node that takes the most time, while the overall computational cost reflects the total effort across all nodes.
@@
As shown in the figure, the proposed and federated methods achieve significantly lower training time thanks to their parallel design. Unlike centralized and swarm learning baselines, they allow simultaneous training across nodes, which improves efficiency and scalability. For computational resource cost, centralized learning shows the best result, as it benefits from a simpler pipeline without requiring additional global classifier training.
5.2 Phase2: Classification on 15 Nodes
Table 3: Phase2 performance comparison on 15-node classification task.